A trader collects the below information to devise his forex management strategy during



the next year:



Spot Rate of USD: INR 74.40



Interest rates in USA: 2.5%



Interest rates in India: 6.5%




a) What is Interest Rate Parity Theory? What will be the expected exchange rate in the above



case, if the interest rate parity theory is assumed to hold good. Also calculate the forward



premium or discount.




b) What is covered interest arbitrage? If there is a one-year forward contract available at



INR 75.60, is there a CIA possible, and if yes, write the steps to earn arbitrage profit (use



an equivalent amount of INR 1000000)

1
Expert's answer
2022-03-01T10:32:32-0500

"IRP=S_{t, \frac{a}{b}}\frac{1+i_a}{1+i_b}"

"= S_{t, \frac{a}{b}}\frac{1+0.25}{1+0.65}"

"=76" %

Covered interest arbitrage is an arbitrage trading strategy where the trader capitalizes on the interest rate differential between two countries by using a forward contract to cover exchange rate risk.